OpenStudy (gabbyalicorn):

A porch has the shape of a semicircle with a diameter of 7 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate π

OpenStudy (gabbyalicorn):

@gaos @Mehek14

OpenStudy (gabbyalicorn):

so im guessing 3.14 x 7?

Mehek (mehek14):

\(\dfrac{πr^2}{2}\)

Mehek (mehek14):

so \(\dfrac{3.14*7^2}{2}\)

Mehek (mehek14):

what is \(7^2\)

OpenStudy (gabbyalicorn):

153.86?

OpenStudy (gabbyalicorn):

oh 7^2 49

Mehek (mehek14):

\(\dfrac{153.86}{2}\)

OpenStudy (gabbyalicorn):

76.93

Mehek (mehek14):

that's the area

OpenStudy (gabbyalicorn):

ty!

OpenStudy (gabbyalicorn):

wait but it says the approximate area so it would be 77? @Mehek14

OpenStudy (gabbyalicorn):

@gaos

OpenStudy (anonymous):

yes it would

OpenStudy (gabbyalicorn):

okay, thanks :)